OpenStudy (anonymous):

factor 1/18x^2-x+9/2

OpenStudy (anonymous):

Note that \(\large \frac{1}{18}x^2-x+\frac{9}{2} = \frac{1}{18}(x^2-18x+81)\). Do you notice anything thing nice about that quadratic function now? :-)

OpenStudy (anonymous):

That looks so much better, fractions scare me. Thank you!